XIV Seminario Nazionale di Fisica Teorica

Seconda settimana

9:00  10:00  11:00  12:00  16:00  16:30  17:30
Mar30 Romanino Romanino  Strumia  Strumia  Romanino
Mer31 Feruglio  Feruglio  Strumia  Romanino  Strumia
Gio 1  Zwirner  Zwirner  Strumia  Romanino  Zwirner
Ven 2 Feruglio  Feruglio  Zaffaroni  Zaffaroni  Zaffaroni  Feruglio
Sab 3  Zwirner  Zwirner  Zaffaroni  Zaffaroni
Dom 4  Giornata libera
Lun 5 Gliozzi Gliozzi De Forcrand De Forcrand De Forcrand Gliozzi
Mar 6 Di Giacomo Di Giacomo De Forcrand De Forcrand  Seminar 1
Mer 7 Di Giacomo Di Giacomo Greensite  Zakharov  Seminar 2  Di Giacomo
Gio 8 Di Giacomo Zakharov  Greensite  Greensite  Zakharov  Greensite
Ven 9 Zakharov  Zakharov  Greensite  Greensite

• Philippe De Forcrand, Static q̄q potentials and algorithms

• Adriano Di Giacomo, Monopole condensation and confinement

• Ferdinando Gliozzi, Introduction to Lattice Gauge Theory

• Jeff Greensite, Central vortices and confinement

• Valentine Zakharov, Extended objects and confinement

Please notice that the Coffe break will be served at 10:50 thru 11:10

Seminari

1. Michele Pepe, (ITP, Bern), Confinement without a center: the exceptional gauge group G(2)

2. Massimo D’Elia, (Phys.Dept, University of Genova), From Confined to Deconfined Matter: The Phase Diagram of QCD